# Paul-Erik Haacker, Remco I. Leine, Renu Chaudhary, Kai Diethelm, André Schmidt, Safoura Hashemishahraki

### v2

@@ 4 Linear Time-Invariant Fractional-Order Systems
In order to demonstrate the difficulties in finding analytical solutions of fractional-order Caputo differential equations with infinite memory, we first consider linear time-invariant differential equations. Recall the classical initial value problem with zero lower bound [Diethelm2010]
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast 0}^{\alpha}}z(t)=Az(t),\quad t\geq 0,\quad z(0)=z_{0}\in\mathbb{R}^{n},$ (30)
@@
yields (in the semisimple case) a complete basis of solutions of (30), where $E_{\alpha}$ denotes the one parameter Mittag-Leffler function and $v_{z}\in\mathbb{R}^{n}$. Its Caputo derivative reads ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast 0}^{\alpha}}z(t)=\lambda z(t)$ [Diethelm2010], yielding the characteristic equation
@@
Solutions $\lambda$ of the characteristic equations (32) are directly related to the stability properties of the trivial solution. This is stated in the following Theorem taken from [Diethelm2010, Theorem 7.20] where the definitions of stability terms are given in the same reference.
@@
where $v_{u}\in\mathbb{R}^{n}$ and $\mathrm{Re}(s)>0$. This is a natural choice, as its Liouville-Weyl derivative reads $^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}u(t)=s^{\alpha}u(t)$ [samko1993fractional], yielding the characteristic equation
@@
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\operatorname{D}_{\ast 0}^{\alpha}}u(t)=Au(t)-\mathcal{F}u_{0}(t),\quad t\geq 0,\quad u(0)=u_{0}(0)$ (39)
with the forcing term $\mathcal{F}u_{0}$ of Definition 3.2. Applying the variation-of-constants formula [Diethelm2010], problem (39) has the solution
@@
where $\tilde{A}=\operatorname{diag}(A,A)$. Clearly, the two components of this system are decoupled, and each of them is equivalent to the original initial value problem, so this formulation might appear unnecessarily complicated. It has the advantage, however, that we can now directly use the methods and results of [Diethelm22], which hold for planar systems.
@@
It follows that $Q(s)=0$ for some $s\in\mathbb{C}$ if and only if $s^{\alpha}=A$. But now, for any $s\in\mathbb{C}$ we have that $\arg(s)\in(-\pi,\pi]$, and hence $\arg(s^{\alpha})=\alpha\cdot\arg(s)\in(-\alpha\pi,\alpha\pi]$. Since in our case $\alpha\in(0,1)$, we thus conclude that $|\arg(s^{\alpha})|\neq\pi$, and so the equation $s^{\alpha}=A$ cannot have a solution. Thus, the characteristic function does not possess any zeros in $\mathbb{C}$, and so in particular it does not have any zeros in the closed right half of the complex plane. This observation allows us to invoke [Diethelm22, Theorem 4] which immediately yields the desired result.
@@
Here we use that, since $E_{\alpha,\alpha}(A(\cdot)^{\alpha})$ is monotonically decreasing whenever $A<0$ (this immediately follows from the complete monotonicity of the function $E_{\alpha,\alpha}(-(\cdot))$ shown in [GorenfloEtAl2020, § 4.9.2]), it holds that
@@ 5 Fractional-Order Linear Time-Periodic Systems
and $y_{0}\in{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X(-\infty,t_{0}]}$. The coordinate $y(t)$ is introduced to describe the local behavior of a small perturbation from the periodic solution $x_{\mathrm{p}}$. Thus, asymptotic stability and instability of the periodic solution is given whenever the equilibrium $y\equiv 0$ of (53) is asymptotically stable and unstable, respectively.
The choice of the lower bound -$\infty$ in the differential equation (12a) is necessary when we want to study periodic solutions $x_{\mathrm{p}}(t)=x_{\mathrm{p}}(t+T)$ for all $t$ and some $T>0$ of (12). A version of (12) with zero lower bound cannot have a periodic solution [Tavazoei2009].
@@
whenever $\operatorname{Re}\lambda\geq 0$ and where $(\cdot)^{\alpha}$ denotes the principal value, $\alpha\in(0,1)$ [samko1993fractional]. One may compute the fractional derivative of the Floquet form (55) by interchanging the order of summation and fractional differentiation, justified by Assumption (A2),
@@
The technical Assumptions (A1) and (A2) on $J$ and $p$, respectively, ensure that the reformulation into the nonlinear eigenvalue problem (75) is valid. Both assumptions are following the approach of [Zhou2002] on first-order systems. For a given function $J$, Assumption (A1) is easily checked. However, as the function $p$ is a priori unknown and is only approximated via truncated Fourier series in the following, Assumption (A2) remains a technical assumption.
A nonlinear eigenvalue problem such as (75) as a characteristic equation is not uncommon. We have seen in Section 4, that a similar but simpler eigenvalue problem occurs for time-invariant systems. Another example of nonlinear eigenvalue problems as characteristic equations shows up in delay differential equations [Stepan2011] for both the time invariant and time periodic cases. Therein, the nonlinearity in the eigenvalue problem is an exponential of the eigenvalue $\lambda$.
The constraint $\operatorname{Re}\lambda\geq 0$ is neccessary, since otherwise the Liouville-Weyl derivative does not exist. This shows, that an extension of Floquet form solutions as in [Traversa2020] to the system (53) is generally not possible. For a certain class of systems with memory terms, results on forms of solutions are available [Traversa2020]. However, fractionally damped mechanical systems [Hinze2020]
@@
just barely escape the requirements of [Traversa2020], since the memory kernel $K(t,\tau)=(t-\tau)^{-\alpha}/\Gamma(1-\alpha)$ does not satisfy $\int_{-\infty}^{t}|K(t,\tau)|\mathrm{d}\tau<\infty$ for any $t$.
@@
By the structure of the fractional Hill matrix (57), given a solution $\lambda$ which makes $\mathbf{H}_{\infty}^{\alpha}(\lambda)$ rank deficient, $\lambda+ik\omega$ for all $k\in\mathbb{Z}$ is also a solution. Hence, the Floquet exponents come in groups, just as in the integer-order case $\alpha=1$ [Bayer2023a]. This comes as no surprise, as the property is a consequence of the ansatz (55).
@@
where $J$ is $T$-periodic. Inspired by the form of solutions of LTI first-order DEs and LTI-FODEs with lower bound zero, taking an Ansatz as (55) but with Mittag-Leffler functions, will greatly complicate setting up a Hill-type problem due to the lack of a multiplicative identity of the Mittag-Leffler function [Diethelm2010]
@@
belongs to the class of Nonlinear Eigenvalue Problems (NEPs) described by [Guettel2017]
@@
where the eigenvalue $z$ may appear nonlinearly, but the corresponding eigenvector $v$ linearly in the equation. Numerical treatment of these problems is extensively discussed in [Guettel2017].
@@
In the case $\alpha=1$, $\lambda$ is a linear eigenvalue of the Hill matrix. Indeed, $\mathbf{H}_{N}^{(1)}(\lambda)=\mathbf{H}_{N}-\lambda I$ where $\mathbf{H}_{N}=\mathbf{H}_{N}^{(1)}(0)$ is the classical Hill-matrix [Bayer2023a].
@@
For the sake of simplicity, consider the scalar case $n=1$. We apply [Guettel2017, Thm 2.19] noting that the fractional Hill matrix
@@
Figure 3: Numerical solutions of Example 5.12 with constant initial condition $y_{0}(t)=1$ for $t\leq 0$, $\alpha=1/2$, a=1a=-1 and varying parameter $b$. We used Garappa’s implementation fde12.m in Matlab of the FracPECE scheme [diethelm1998fracpece].
@@
Figure 6: Results of numerical time integration $y_{\mathrm{sim}}(t)$ of the system (83) with $\alpha=1/2$, a=1a=-1, $N=10$ and $b=2.2$, compared with the solution obtained by the Hill method $y_{\mathrm{Hill}}(t)$. We used the numerical scheme of [Chaudhary2024] implemented by Afshin Farhadi for problems of the form $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x=f(t,x)$ to initial condition $x(0)=x_{0}$. We implemented the infinite-memory initial condition $y_{0}$ as a forcing term on the right-hand side.
@@ 6 Conclusion and Open Questions
    In the case of LTP-FODEs, how can one construct a monodromy operator and is it possible to find Floquet multipliers? Since the history of trajectories is a part of FODEs, we conjecture that a monodromy operator $\Phi_{T}:X(-\infty,0)\to X(-\infty,T)$ is required for the study of LTP systems analogous to the case of functional differential equations [Hale1977]. The spectrum of this operator will map the stability properties of the respective system. Consequently, the study of fundamental solution operators for LTP-FODEs is a neccessary first step.
@@
system LTI LTP
order $\alpha$ 1 $(0,1)$ 1 $(0,1)$
@@
EVP $Av=\rho v$ $Av=\rho^{\alpha}v$ $\mathbf{H}_{\infty}v=\rho v$ - $\mathbf{H}_{\infty}^{(\alpha)}(\rho)v=0$
$\operatorname{dom}\rho$ $\rho\in\mathbb{C}$ $\operatorname{Re}\rho\geq 0$ $\rho\in\mathbb{C}$ - $\operatorname{Re}\rho\geq 0$
@@ Acknowledgements
Assume w.l.o.g. $t_{0}=0$. We first show that that $\mathcal{F}x_{0}\in C[0,\infty)$. In other words, we must demonstrate that, for all $t\in[0,\infty)$, $\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t+\varepsilon)\to\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t)$ as $\varepsilon\to 0$.
@@
$\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t+\varepsilon)-\mathcal{F}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(t)=\frac{1}{\Gamma(1-\alpha)}(B_{1}+B_{2})$ (87)
@@
$B_{1}$ $=\int_{-\eta}^{0}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)\,\mathrm{d}\tau$ (88)
@@
$B_{2}$ $=\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)\,\mathrm{d}\tau$ (89)
where $\eta>0$ is the fixed number from the property that ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}\in L^{\infty}[-\eta,0]$. Recalling that $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}$$^{\prime}$ is bounded on $\eta,0]$, we see that
$|B_{1}|$ $\leq\sup_{\tau\in[-\eta,0]}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)|\cdot\int_{-\eta}^{0}\left|(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right|\,\mathrm{d}\tau$ (90)
$=\sup_{\tau\in[-\eta,0]}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}^{\prime}(\tau)|\frac{1}{1-\alpha}\left[(t+\eta)^{1-\alpha}-(t+\varepsilon+\eta)^{1-\alpha}-t^{1-\alpha}+(t+\varepsilon)^{1-\alpha}\right],$ (91)
@@
$B_{21}$ $=\left.\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]^{\vphantom{b}}\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)\right|_{\tau=-\infty}^{\tau=-\eta}$ (93)
@@
$B_{22}$ $=-\alpha\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right]{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)\,\mathrm{d}\tau.$
Clearly, since $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}$} is bounded and the term in square brackets in (93) tends to zero as $\tau\to$-$\infty$, we find that
$B_{21}=\left[(t+\varepsilon+\eta)^{-\alpha}-(t+\eta)^{-\alpha}\right]\cdot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(-\eta)\to 0$ (94)
for $\varepsilon\to 0$ since $t\geq 0$ and $\eta>0$. Finally, once again using the boundedness of $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}$} and applying the mean value theorem of differential calculus to the function $(\cdot)^{-1-\alpha}$ on the interval $t-\tau,t+\varepsilon-\tau]$, we see that
$|B_{22}|$ $\leq\alpha\sup_{\tau\leq-\eta}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)|\int_{-\infty}^{-\eta}\left|(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right|\,\mathrm{d}\tau$ (95)
$=\alpha(\alpha+1)\sup_{\tau\leq-\eta}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)|\cdot\varepsilon\int_{-\infty}^{-\eta}\left|\xi_{\tau,\varepsilon}\right|^{-2-\alpha}\,\mathrm{d}\tau$ (96)
@@
$|B_{22}|$ $\leq\varepsilon~\alpha(\alpha+1)\sup_{\tau\leq-\eta}|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{0}}(\tau)|\cdot\int_{-\infty}^{-\eta}(t-\tau)^{-2-\alpha}\,\mathrm{d}\tau$ (97)

### v3

@@ 4 Linear Time-Invariant Fractional-Order Systems
In order to demonstrate the difficulties in finding analytical solutions of fractional-order Caputo differential equations with infinite memory, we first consider linear time-invariant differential equations. Recall the classical initial value problem with zero lower bound [10]
${\operatorname{D}_{\ast 0}^{\alpha}}z(t)=Az(t),\quad t\geq 0,\quad z(0)=z_{0}\in\mathbb{R}^{n},$ (30)
@@
yields (in the semisimple case) a complete basis of solutions of (30), where $E_{\alpha}$ denotes the one parameter Mittag-Leffler function and $v_{z}\in\mathbb{R}^{n}$. Its Caputo derivative reads ${\operatorname{D}_{\ast 0}^{\alpha}}z(t)=\lambda z(t)$ [10], yielding the characteristic equation
@@
Solutions $\lambda$ of the characteristic equations (32) are directly related to the stability properties of the trivial solution. This is stated in the following Theorem taken from [10, Theorem 7.20] where the definitions of stability terms are given in the same reference.
@@
where $v_{u}\in\mathbb{R}^{n}$ and $\mathrm{Re}(s)>0$. This is a natural choice, as its Liouville-Weyl derivative reads $^{\mathrm{C}}\operatorname{D}_{-\infty}^{\alpha}u(t)=s^{\alpha}u(t)$ [31], yielding the characteristic equation
@@
${\operatorname{D}_{\ast 0}^{\alpha}}u(t)=Au(t)-\mathcal{F}u_{0}(t),\quad t\geq 0,\quad u(0)=u_{0}(0)$ (39)
with the forcing term $\mathcal{F}u_{0}$ of Definition 3.2. Applying the variation-of-constants formula [10], problem (39) has the solution
@@
where $\tilde{A}=\operatorname{diag}(A,A)$. Clearly, the two components of this system are decoupled, and each of them is equivalent to the original initial value problem, so this formulation might appear unnecessarily complicated. It has the advantage, however, that we can now directly use the methods and results of [9], which hold for planar systems.
@@
It follows that $Q(s)=0$ for some $s\in\mathbb{C}$ if and only if $s^{\alpha}=A$. But now, for any $s\in\mathbb{C}$ we have that $\arg(s)\in(-\pi,\pi]$, and hence $\arg(s^{\alpha})=\alpha\cdot\arg(s)\in(-\alpha\pi,\alpha\pi]$. Since in our case $\alpha\in(0,1)$, we thus conclude that $|\arg(s^{\alpha})|\neq\pi$, and so the equation $s^{\alpha}=A$ cannot have a solution. Thus, the characteristic function does not possess any zeros in $\mathbb{C}$, and so in particular it does not have any zeros in the closed right half of the complex plane. This observation allows us to invoke [9, Theorem 4] which immediately yields the desired result.
@@
Here we use that, since $E_{\alpha,\alpha}(A(\cdot)^{\alpha})$ is monotonically decreasing whenever $A<0$ (this immediately follows from the complete monotonicity of the function $E_{\alpha,\alpha}(-(\cdot))$ shown in [13, § 4.9.2]), it holds that
@@ 5 Fractional-Order Linear Time-Periodic Systems
and $y_{0}\in{X(-\infty,t_{0}]}$. The coordinate $y(t)$ is introduced to describe the local behavior of a small perturbation from the periodic solution $x_{\mathrm{p}}$. Thus, asymptotic stability and instability of the periodic solution is given whenever the equilibrium $y\equiv 0$ of (53) is asymptotically stable and unstable, respectively.
The choice of the lower bound -$\infty$ in the differential equation (12a) is necessary when we want to study periodic solutions $x_{\mathrm{p}}(t)=x_{\mathrm{p}}(t+T)$ for all $t$ and some $T>0$ of (12). A version of (12) with zero lower bound cannot have a periodic solution [34].
@@
whenever $\operatorname{Re}\lambda\geq 0$ and where $(\cdot)^{\alpha}$ denotes the principal value, $\alpha\in(0,1)$ [31]. One may compute the fractional derivative of the Floquet form (55) by interchanging the order of summation and fractional differentiation, justified by Assumption (A2),
@@
The technical Assumptions (A1) and (A2) on $J$ and $p$, respectively, ensure that the reformulation into the nonlinear eigenvalue problem (75) is valid. Both assumptions are following the approach of [38] on first-order systems. For a given function $J$, Assumption (A1) is easily checked. However, as the function $p$ is a priori unknown and is only approximated via truncated Fourier series in the following, Assumption (A2) remains a technical assumption.
A nonlinear eigenvalue problem such as (75) as a characteristic equation is not uncommon. We have seen in Section 4, that a similar but simpler eigenvalue problem occurs for time-invariant systems. Another example of nonlinear eigenvalue problems as characteristic equations shows up in delay differential equations [22] for both the time invariant and time periodic cases. Therein, the nonlinearity in the eigenvalue problem is an exponential of the eigenvalue $\lambda$.
The constraint $\operatorname{Re}\lambda\geq 0$ is neccessary, since otherwise the Liouville-Weyl derivative does not exist. This shows, that an extension of Floquet form solutions as in [35] to the system (53) is generally not possible. For a certain class of systems with memory terms, results on forms of solutions are available [35]. However, fractionally damped mechanical systems [18]
@@
just barely escape the requirements of [35], since the memory kernel $K(t,\tau)=(t-\tau)^{-\alpha}/\Gamma(1-\alpha)$ does not satisfy $\int_{-\infty}^{t}|K(t,\tau)|\mathrm{d}\tau<\infty$ for any $t$.
@@
By the structure of the fractional Hill matrix (57), given a solution $\lambda$ which makes $\mathbf{H}_{\infty}^{\alpha}(\lambda)$ rank deficient, $\lambda+ik\omega$ for all $k\in\mathbb{Z}$ is also a solution. Hence, the Floquet exponents come in groups, just as in the integer-order case $\alpha=1$ [2]. This comes as no surprise, as the property is a consequence of the ansatz (55).
@@
where $J$ is $T$-periodic. Inspired by the form of solutions of LTI first-order DEs and LTI-FODEs with lower bound zero, taking an Ansatz as (55) but with Mittag-Leffler functions, will greatly complicate setting up a Hill-type problem due to the lack of a multiplicative identity of the Mittag-Leffler function [10]
@@
belongs to the class of Nonlinear Eigenvalue Problems (NEPs) described by [14]
@@
where the eigenvalue $z$ may appear nonlinearly, but the corresponding eigenvector $v$ linearly in the equation. Numerical treatment of these problems is extensively discussed in [14].
@@
In the case $\alpha=1$, $\lambda$ is a linear eigenvalue of the Hill matrix. Indeed, $\mathbf{H}_{N}^{(1)}(\lambda)=\mathbf{H}_{N}-\lambda I$ where $\mathbf{H}_{N}=\mathbf{H}_{N}^{(1)}(0)$ is the classical Hill-matrix [2].
@@
For the sake of simplicity, consider the scalar case $n=1$. We apply [14, Thm 2.19] noting that the fractional Hill matrix
@@
Figure 3: Numerical solutions of Example 5.12 with constant initial condition $y_{0}(t)=1$ for $t\leq 0$, $\alpha=1/2$, a=1a=-1 and varying parameter $b$. We used Garappa’s implementation fde12.m in Matlab of the FracPECE scheme [8].
@@
Figure 6: Results of numerical time integration $y_{\mathrm{sim}}(t)$ of the system (83) with $\alpha=1/2$, a=1a=-1, $N=10$ and $b=2.2$, compared with the solution obtained by the Hill method $y_{\mathrm{Hill}}(t)$. We used the numerical scheme of [4] implemented by Afshin Farhadi for problems of the form $^{\mathrm{C}}\operatorname{D}_{0}^{\alpha}x=f(t,x)$ to initial condition $x(0)=x_{0}$. We implemented the infinite-memory initial condition $y_{0}$ as a forcing term on the right-hand side.
@@ 6 Conclusion and Open Questions
    In the case of LTP-FODEs, how can one construct a monodromy operator and is it possible to find Floquet multipliers? Since the history of trajectories is a part of FODEs, we conjecture that a monodromy operator $\Phi_{T}:X(-\infty,0)\to X(-\infty,T)$ is required for the study of LTP systems analogous to the case of functional differential equations [16]. The spectrum of this operator will map the stability properties of the respective system. Consequently, the study of fundamental solution operators for LTP-FODEs is a neccessary first step.
@@
System LTI LTI LTI LTP LTP LTP
order $\alpha$ 1 $(0,1)$ $(0,1)$ 1 $(0,1)$ $(0,1)$
@@
EVP $Av=\rho v$ $Av=\rho v$ $Av=\rho^{\alpha}v$ $\mathbf{H}_{\infty}v=\rho v$ - $\mathbf{H}_{\infty}^{(\alpha)}(\rho)v=0$
$\operatorname{dom}\rho$ $\rho\in\mathbb{C}$ $\rho\in\mathbb{C}$ $\operatorname{Re}\rho\geq 0$ $\rho\in\mathbb{C}$ - $\operatorname{Re}\rho\geq 0$
@@ Acknowledgements
Assume w.l.o.g. $t_{0}=0$. We first show that that $\mathcal{F}x_{0}\in C[0,\infty)$. In other words, we must demonstrate that, for all $t\in[0,\infty)$, $\mathcal{F}{x_{0}}(t+\varepsilon)\to\mathcal{F}{x_{0}}(t)$ as $\varepsilon\to 0$.
@@
$\mathcal{F}{x_{0}}(t+\varepsilon)-\mathcal{F}{x_{0}}(t)=\frac{1}{\Gamma(1-\alpha)}(B_{1}+B_{2})$ (87)
@@
$B_{1}$ $=\int_{-\eta}^{0}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]{x_{0}}^{\prime}(\tau)\,\mathrm{d}\tau$ (88)
@@
$B_{2}$ $=\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]{x_{0}}^{\prime}(\tau)\,\mathrm{d}\tau$ (89)
where $\eta>0$ is the fixed number from the property that ${x_{0}}^{\prime}\in L^{\infty}[-\eta,0]$. Recalling that $x_{0}$$^{\prime}$ is bounded on $\eta,0]$, we see that
$|B_{1}|$ $\leq\sup_{\tau\in[-\eta,0]}|{x_{0}}^{\prime}(\tau)|\cdot\int_{-\eta}^{0}\left|(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right|\,\mathrm{d}\tau$ (90)
$=\sup_{\tau\in[-\eta,0]}|{x_{0}}^{\prime}(\tau)|\frac{1}{1-\alpha}\left[(t+\eta)^{1-\alpha}-(t+\varepsilon+\eta)^{1-\alpha}-t^{1-\alpha}+(t+\varepsilon)^{1-\alpha}\right],$ (91)
@@
$B_{21}$ $=\left.\left[(t+\varepsilon-\tau)^{-\alpha}-(t-\tau)^{-\alpha}\right]^{\vphantom{b}}\cdot{x_{0}}(\tau)\right|_{\tau=-\infty}^{\tau=-\eta}$ (93)
@@
$B_{22}$ $=-\alpha\int_{-\infty}^{-\eta}\left[(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right]{x_{0}}(\tau)\,\mathrm{d}\tau.$
Clearly, since $x_{0}$} is bounded and the term in square brackets in (93) tends to zero as $\tau\to$-$\infty$, we find that
$B_{21}=\left[(t+\varepsilon+\eta)^{-\alpha}-(t+\eta)^{-\alpha}\right]\cdot{x_{0}}(-\eta)\to 0$ (94)
for $\varepsilon\to 0$ since $t\geq 0$ and $\eta>0$. Finally, once again using the boundedness of $x_{0}$} and applying the mean value theorem of differential calculus to the function $(\cdot)^{-1-\alpha}$ on the interval $t-\tau,t+\varepsilon-\tau]$, we see that
$|B_{22}|$ $\leq\alpha\sup_{\tau\leq-\eta}|{x_{0}}(\tau)|\int_{-\infty}^{-\eta}\left|(t+\varepsilon-\tau)^{-1-\alpha}-(t-\tau)^{-1-\alpha}\right|\,\mathrm{d}\tau$ (95)
$=\alpha(\alpha+1)\sup_{\tau\leq-\eta}|{x_{0}}(\tau)|\cdot\varepsilon\int_{-\infty}^{-\eta}\left|\xi_{\tau,\varepsilon}\right|^{-2-\alpha}\,\mathrm{d}\tau$ (96)
@@
$|B_{22}|$ $\leq\varepsilon~\alpha(\alpha+1)\sup_{\tau\leq-\eta}|{x_{0}}(\tau)|\cdot\int_{-\infty}^{-\eta}(t-\tau)^{-2-\alpha}\,\mathrm{d}\tau$ (97)